A partial relative frequency distribution is given. a. What is the relative frequency of class D (to 2 decimals)? Class Relative Frequency 0.22 0.26 0.09 b. The total sample size is 900. What is the frequency of class D? c. Show the frequency distribution. Class Frequency Total 900 d. Show the percent frequency distribution (to whole number) Class Percent Frequency Total 100%
Answer #1

Answer

(A) we know that the sum of relative frequency must be equal to 1

So, we can write it as

f(A) + f(B) + f(C) + f(D) = 1

using the data table above, putting the values of frequencies, we get

0.22+0.26+0.09 + f(D) = 1

0.53+f(D) = 1

subtracting 0.53 on each side, we get

f(D) = 1 - 0.57 = 0.43

So, relative frequency of D is 0.43

(b) total sample size = 900

relative frequency of D is 0.43

So, Frequency of class D = relative frequency*sample size = 0.43*900 = 387

(c)

Class Frequency
A relative frequency*sample size= 0.22*900=198
B relative frequency*sample size= 0.26*900=234
C relative frequency*sample size= 0.09*900=81
D relative frequency*sample size= 0.43*900=387

(D) Percent frequency = relative frequency*100

Class Percent Frequency
A 0.22*100 = 22%
B 0.26*100 = 26%
C 0.09*100 = 9%
D 0.43*100 = 43%